Circular No. 5860

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993X IN NGC 2276 A. V. Filippenko and T. Matheson, University of California at Berke-

ley, report: “CCD spectrograms (range 310-990 nm, resolution 0.7 nm) obtained on Sept. 10 UT with the Shane 3-m reflector at Lick Observatory reveal that SN 1993X (cf. IAUC 5850, 5853) is of type II. The spectrum is somewhat peculiar, however, being dominated by a broad Hα emission line with little or no P-Cyg absorption. A much narrower, but slightly resolved, Hα emission line is also present. These characteristics resemble those of supernovae whose ejecta experience considerable interaction with circumstellar gas. Observations at radio wavelengths would be worthwhile. The recession velocity of the parent galaxy is ∼ 2400 km/s.”

GRO J0422+32 P. Zhao, P. Callanan, M. Garcia, and J. McClintock, Smithsonian As-

trophysical Observatory, write: “CCD photometry obtained at the F. L. Whipple Observatory 1.2-m telescope gives the following R magnitudes: Sept. 8.37 UT, 18.1 ± 0.5; 9.48, 17.7 ± 0.2. Thus, GRO J0422+32 has dimmed by ∼ 2-3 mag since its unexpected brightening on Aug. 10-16 (cf. IAUC 5842, 5844). This may help explain the faintness of the x-ray source observed on Aug. 26 UT (IAUC 5851). Since the optical counterpart’s rapid decline in Mar.-Apr., we have obtained photometry on 16 nights during Apr. 13-28 (V, R), July 18-29 (R), and Sept. 8-9 (R); all these low-state data are consistent with V = 17.6 ± 0.3 and R = 17.8 ± 0.3. We have monitored V and R on 58 nights since 1992 Nov. 18.”

Filippenko and Matheson also report that GRO J0422+32 has faded very rapidly since its second outburst in August. Rough estimates made on Sept. 11 UT with the Shane 3-m reflector yield V = 18.9 ± 0.5, as compared with V = 15.3 ± 0.1 on Aug. 24 (IAUC 5851). The spectrum (range 310-800 nm), although noisy, is reasonably featureless except for prominent, double-peaked hydrogen Balmer emission lines and weaker helium lines.

COMET MUELLER (1993a) Total visual magnitude estimates (cf. IAUC 5838): Aug. 18.99 UT,

11.6 (B. H. Granslo, Fjellhamar, Norway, 0.20-m reflector); 21.45, 11.7 (A. Hale, Orogrande, NM, 0.20-m reflector); 29.19, 10.7 (A. Pereira, Cabo da Roca, Portugal, 0.15-m reflector); Sept. 11.08, 11.0 (Granslo).

1993 September 11 Daniel W. E. Green